OpenStudy (anonymous):

givrn: y=2x^2 -4x+3 find the vertex

OpenStudy (anonymous):

use \[-\frac{b}{2a}\]

OpenStudy (anonymous):

in your example a = 2, b = -4 and \[-\frac{-4}{2\times 2}=1\]

OpenStudy (anonymous):

first coordinate of the vertex is 1, second is what y ou get when you replace x by 1, 2-4+3=1 so vertex is (1,1)

OpenStudy (anonymous):

what would be the intercept

OpenStudy (anonymous):

there is no x intercept since this is a parabola that faces up and the vertex is above the x - axis

OpenStudy (anonymous):

y intercept

OpenStudy (anonymous):

the y-intercept is what you get when you replace x by 0, which is the constant . in this case it is 3, so y-intercept is (0,3)